A ray of light in corn oil is incident striking with an angle of incidence of 40 degrees on water. Calculate the angle of refraction to the nearest degree


1
Expert's answer
2021-07-01T12:50:35-0400

Given:

"n_{\rm oil}=1.48"

"n_{\rm water}=1.33"

"\theta_i=40^{\circ}"

The Snell's law says

"n_{\rm oil}\sin\theta_i=n_{\rm water}\sin\theta_r"

Hence,

"\sin\theta_r=n_{\rm oil}\sin\theta_i/n_{\rm water}\\
=1.48\sin 40^{\circ}/1.33=0.715"

"\theta_r=46^{\circ}"